Find the equation of a line that is parallel to x + 3y = 2 and containing the point (0, 0).
Answer by mananth(16946)

1 x + 3 y = 2
Find the slope of this line
make y the subject
3 y = -1 x + 2
Divide by 3
y = -0.33 x + 0.67
Compare this equation with y=mx+b
slope m = -0.33 0.8
The slope of a line parallel to the above line will be the same
The slope of the required line will be -0.33
m= -0.33 ,point ( 0 , 0 )
Find b by plugging the values of m & the point in
y=mx+b
0 = 0.00 + b
b= 0
m= -0.33
Plug value of the slope and b in y = mx +b
The required equation isy = -0.33 x + 0
y = - 1/ 3 x + 0
m.ananth@hotmail.ca